OpenStudy (anonymous):

A person spends 30% of his salary on house rent and 55% of the remaining money on food and clothes. Hesaves the remaining amount. If he saves ?59724 yearly, how much amount does he spend on food and clothes per month?

OpenStudy (shamim):

let his salary is \[x\]

OpenStudy (shamim):

he spent for house rent and food, clothing is\[30+55=85\]

OpenStudy (shamim):

so his savings is\[100-85=15percent\]

OpenStudy (shamim):

so\[\frac{ 15 }{ 100 }x=59724\]

OpenStudy (shamim):

try to find \[x\]

OpenStudy (shamim):

his clothing cost wil b\[55 percent \times x\]

OpenStudy (shamim):

\[\frac{ 55 }{ 100 }x\]

OpenStudy (shamim):

u need to divide this amount by 12

OpenStudy (phi):

I would interpret **A person spends 30% of his salary on house rent and 55% of the remaining money on food and clothes.** to mean x = salary 0.3x on rent leaving 0.7x remaining 0.55* 0.7 x on food and clothes If he saves 59724 yearly, you would say x - (0.3x + 0.55*0.7x) = 59724 which says: total salary minus expenses is saved, equal to 59724 you can solve this by simplifying. first, do 0.55*0.7: x - (0.3x + .385x) = 59724 add 0.3x and 0.385 x to get x - 0.685x = 59724 now simplify 1x -0.685 x 0.315x = 59724 divide both sides by 0.315 to get x= 59724/0.315

OpenStudy (anonymous):

59724/12= 4977 (which is 15% of his monthly earnings) so that multiplied by 2 is 9954:which is how much he spends of his house (30%) then 4977/15*55= 18249:which is how much he spends on food and clothing (55%)

OpenStudy (phi):

to finish: total salary for the year is x= 59724/0.315 = 189600 0.55* 0.7 * x on food and clothes per year, so 0.55*0.7*189600 = 72996 food and clothes per year 72996/12 = 6083 food and clothes per month